# Large deviations of current for the symmetric simple exclusion process on a semi-infinite line and on an infinite line with a slow bond

### v2

@@ Abstract
Two of the most influential exact results in classical one-dimensional diffusive transport are about current statistics for the symmetric simple exclusion process in the stationary state on a finite line coupled with two unequal reservoirs at the boundary, and in the non-stationary state on an infinite line. We present the corresponding result for the intermediate geometry of a semi-infinite line coupled with a single reservoir. This result is obtained using the fluctuating hydrodynamics approach of macroscopic fluctuation theory and confirmed by rare event simulations using a cloning algorithm. Our exact result enables us to address the corresponding problem on an infinite line in the presence of a slow bond and several related problems.
large deviations; symmetric simple exclusion process; slow bond
@@
Current fluctuations in non-equilibrium transport have long been a subject of interest in both classical and quantum contexts [1, 2, *2006_Bertini_Non, 4, 5, 6, 7, 8]. The major interest in these investigations lies in the full counting statistics in terms of large deviations of current. Apart from being a contender for extension of free energy for non-equilibrium systems, large deviation function (ldf) in general can characterize various peculiarities of non-equilibrium conditions, such as non-local response, emergent symmetries, and low-dimensional phase transitions [9, 4, 10, 11, 12, 13, 14, 15]. However, estimating ldf poses a major challenge, often requiring specialized integrability techniques [16, 5, 17] tailored to curated models or clever numerical sampling schemes of rare-probability events [18, 19, 20, 21, 22]. Understandably, exact results about ldf play an important role in the landscape of non-equilibrium physics, providing a benchmark for affirming qualitative predictions of approximate methods. Our work in this Letter presents a non-trivial addition to this list of exact results, fostering solutions to certain relevant problems [23, 24, 25, *2016_Franco_Phase].
@@
Among the widely studied stochastic models of classical transport, is the symmetric simple exclusion process (SSEP) [27, 28, 29, 4]. The SSEP, along with its driven variants, has attained the status of a paradigmatic model in non-equilibrium statistical physics [30, 17]. Two celebrated exact results for the SSEP concern the large deviations of current on a finite lattice coupled with two unequal reservoirs, and on an infinite lattice starting with a non-stationary state. The two geometries represent distinct non-equilibrium scenarios: a stationary state for finite systems, and a time-dependent state relaxing towards an asymptotic equilibrium state for infinite systems. They illuminate crucial differences in their fluctuations. The finite system, in the long run, holds no memory of the initial state, while the infinite system exhibits an unusual dependence on the initial state even at long times [31, 32].
@@
The large deviations of current in these two geometries were obtained using the additivity conjecture [33, 34, 35] and integrability methods, such as the diagonalization of tilted matrix [36] or the matrix product states [37, 5] for the finite lattice, and the Bethe ansatz [38] for the infinite lattice. These microscopic results were subsequently verified [2, *2006_Bertini_Non, 39, *2024_Mallick_Exact, 41] using a fluctuating hydrodynamics framework [38, 42, 43]. This framework, now famously known as the macroscopic fluctuation theory (MFT) [44, *2002_Bertini_Macroscopic, 2, *2006_Bertini_Non], emerged in the early 2000s from the seminal works of Bertini, De Sole, Gabrielli, Jona-Lasinio, and Landim, which presented a general approach for characterizing non-equilibrium fluctuations of diffusive systems. MFT has successfully led to exact results of large deviations in exclusion processes [2, *2006_Bertini_Non, 46, *2008_Tailleur_Mapping, 42, 39, 41] and related transport models [42, 48, *2022_Bettelheim_Full, 50, 51].
@@
In this Letter, we consider the intermediate scenario: the SSEP on a semi-infinite lattice [52, 53] coupled to a boundary reservoir with a density that is different from the initial bulk density of the system. This elucidates a non-equilibrium regime evolving into an asymptotic state in equilibrium with the reservoir, although the dynamical quantities remain sensitive to the initial conditions [54]. Only limited results [55, 56, 57, 54] are known for this intermediate geometry, particularly because extending the aforementioned integrability methods for this geometry proves challenging. Even a solution via MFT has remained elusive.
@@
We overcame these difficulties of the semi-infinite-problem by uncovering a novel mapping to the infinite-line problem that led to an exact result for the ldf of current for SSEP on the semi-infinite geometry, which is our main result. Additionally, this exact result enables us to characterize current fluctuations in the presence of a slow bond, both in the semi-infinite and infinite geometries. The mapping to the infinite-problem is possible due to certain symmetries of the Euler-Lagrange equations and the boundary conditions of the associated variational problem within MFT. These techniques, however, could not be obviously exploited within microscopic approaches. Our exact result is a testament to the power of MFT for non-equilibrium systems that are otherwise formidable to approach.
@@
Model and main result: The SSEP on a semi-infinite lattice (see Fig. 1), with sites indexed by $i=1,2,\cdots$, is composed of continuous-time hard-core random walkers hopping to adjacent sites with a unit rate provided the target site is empty. At the boundary site $i=1$, particles are injected following exclusion with rate $\gamma\rho_{a}$ and removed at rate $\gamma(1-\rho_{a})$, which models [4] coupling with a reservoir of density $\rho_{a}$. For each site $i$ at a given time $\tau$, we assign a binary occupation number $n_{i}(\tau)$ that takes values 00 and $1$, depending on whether the site is empty or occupied, respectively. Initially, sites are filled following Bernoulli distribution with a uniform average density $\rho_{b}$.
@@
Figure 1: The SSEP on a semi-infinite lattice coupled with a reservoir of density $\rho_{a}$ with a coupling strength $\gamma>0$.
@@
Our main result concerns the time-integrated current, $Q_{T}$, which represents the total flux of particles from the reservoir into the system over a time-period $T$. In the hydrodynamic description [58], expressed in terms of coarse-grained density $\rho(i/\sqrt{T},\tau/T)\simeq n_{i}(\tau)$, the flux
@@
$\big{<}\mathrm{e}^{\lambda Q_{T}}\big{>}\sim\mathrm{e}^{\sqrt{T}\mu_{\text{si}}(\lambda,\rho_{a},\rho_{b})}.$ (2a)
Here, the scaled cumulant generating function (scgf) $\mu_{\text{si}}(\lambda,\rho_{a},\rho_{b})$ (subscript ‘si’ denotes semi-infinite) is given by
$\mu_{\text{si}}(\lambda,\rho_{a},\rho_{b})=\int_{-\infty}^{\infty}\frac{\mathrm{d}k}{2\pi}\log{\big{[}1+4\,\omega\,(1+\omega)\,\mathrm{e}^{-k^{2}}\big{]}}$ (2b)
where $\omega$ is a function of the parameters defined as [38, 32]
$\omega(\lambda,\rho_{a},\rho_{b})=\rho_{a}(1-\rho_{b})(\mathrm{e}^{\lambda}-1)+\rho_{b}(1-\rho_{a})(\mathrm{e}^{-\lambda}-1).$ (2c)
This dependence of scgf on $\lambda,\rho_{a}$, and $\rho_{b}$ through a single function $\omega$ arises from a symmetry of the underlying dynamics of the SSEP [36, 38, 59]. The expression (2b) is consistent with our earlier [54] result in the low-density limit. It is instructive to compare (2b) with the corresponding result for the infinite line [38].
The scgf (2b) encapsulates all cumulants of $Q_{T}$ in the large $T$ limit. While the first three cumulants were initially reported in [54], we present here the fourth cumulant for $\rho_{a}=\rho_{b}=\rho$
Similar to the infinite line case [31], the scgf (2b) also admits the Gallavotti-Cohen-type fluctuation symmetry
@@
Remarkably, the expression (2b) admits a Fredholm determinant representation
$\mu_{\text{si}}(\lambda,\rho_{a},\rho_{b})=\frac{1}{2\pi}\log{\det{[\mathbbm{1}+4\omega(1+\omega)K(x,y)]}},$ (5)
with the kernel $K(x,y)=\delta(x-y)\exp{(\partial_{x}^{2})}$. Analogous Fredholm determinant structure is observed in related contexts for the current distribution in the TASEP [60] and the ASEP [61, 62]. This representation underscores a connection between TASEP and random matrix theory, as emphasized by Johansson [60]. However, a similar relationship for the SSEP is not apparent. Interestingly, an alternative representation of (2b),
$\mu_{\text{si}}(\lambda,\rho_{a},\rho_{b})=\frac{1}{2\sqrt{\pi}}\sum_{n=1}^{\infty}\frac{(-1)^{n+1}}{n^{3/2}}\,[4\,\omega\,(1+\omega)]^{n}$ (6)
bears a striking resemblance  [^1] This perplexing similarity was brought to our attention by Gunter M. Schütz. to the scgf for the number of surviving particles in an assembly of annihilating random walkers (see eq. (9.106) in [16]).
Figure 2: Solid red line represents the theoretical result of scgf (2b) for $\rho_{a(b)}=0.5$, with blue dots representing the corresponding simulation result obtained by the cloning algorithm. The results closely match in the range $-1.5\leq\lambda\leq 1.5$, with the deviations shown in the inset. For comparison, the corresponding scgf for infinite line $\mu_{\text{inf}}$ [38] is shown in black dashed line.
Numerical verification: For confirming our result (2b) based on the hydrodynamic description of SSEP, we have independently generated the scgf using numerical simulation based on a continuous-time cloning algorithm [58, 18, 19] with $10^{5}$ clones and measured over duration $T=250$. The simulation result, plotted in Fig. 2, shows a good agreement with our theoretical result (2b) for a reasonably large value of $\lambda$. Deviations emerging at larger values of $\lambda$ are a consequence of finite-size effects [58].
LDF: The scaling in (2a) corresponds to a large deviation asymptotics of the probability
@@
where $\phi_{\text{si}}(j)$ is the ldf, related to $\mu_{\text{si}}(\lambda)$ (reference to $\rho_{a(b)}$ is ignored) by a Legendre-Fenchel transformation [12] $\phi_{\text{si}}(j)=\max_{\lambda}{[j\lambda-\mu_{\text{si}}(\lambda)]}$. Using the transformation, it is immediate that (4) reflects the symmetry
@@
Similarly, the asymptotics $\mu_{\text{si}}(\lambda,\rho_{a},\rho_{b})\simeq\frac{4\sqrt{2}}{3\pi}\,{(\log{\omega})}^{3/2}$ of (2b) for large positive $\lambda$ correspond [38] to the asymptotics of the ldf
@@
The symmetry relation (8) and the asymptotics (9) are confirmed in the plot of the ldf in Fig. 3, generated by numerically evaluating the Legendre-Fenchel transformation of the scgf (2b).
Figure 3: The ldf indicated in solid red line, with the super-Gaussian asymptotes (9) for large $|j|$ indicated in blue dot-dashed lines. The black dashed line denotes the Gaussian ldf with the appropriate mean $\langle Q_{T}\rangle$ and variance $\langle Q_{T}^{2}\rangle_{c}$. The plots are for $\rho_{a}=0.6$ and $\rho_{b}=0.3$. The inset confirms the symmetry (8).
Derivation: In the following, we outline our derivation of (2b) within the fluctuating hydrodynamics framework of MFT. The crucial idea behind MFT is to recognize the relevant hydrodynamic modes for a coarse-grained description of the dynamics and characterize the probability of their evolution in terms of an Action, which is analogous to the Martin-Siggia-Rose-Janssen-De Dominicis (MSRJD) Action [64, 65, 66, 67] of the associated fluctuating hydrodynamics equation. For SSEP, the relevant hydrodynamic mode is the locally conserved density $\rho(x,t)$ evolving by [54, 58]
@@
where $\sigma(\rho)=2\rho(1-\rho)$, and $\eta(x,t)$ is a delta-correlated Gaussian white noise with unit covariance. Corresponding MSRJD-Action on the semi-infinite line is [54, 58]
$S_{0}=\mathcal{F}+\int_{0}^{1}\mathrm{d}t\int_{0}^{\infty}\mathrm{d}x\bigg{[}\widehat{\rho}\,\partial_{t}\rho-\frac{\sigma(\rho)\,(\partial_{x}\widehat{\rho})^{2}}{2}\,+\partial_{x}\rho\,\partial_{x}\widehat{\rho}\bigg{]}$ (11)
where $\widehat{\rho}$ is the response field and $\mathcal{F}$ incorporates contributions from fluctuations in the initial state [31, 54]. Within this description, the generating function $\langle\mathrm{e}^{\lambda Q_{T}}\rangle=\int\mathcal{D}\left[\rho,\widehat{\rho}\right]\,\mathrm{e}^{-\sqrt{T}S_{\lambda}\left[\rho,\widehat{\rho}\right]}$ with $S_{\lambda}=S_{0}-\lambda Q_{T}/\sqrt{T}$ and $Q_{T}$ in (1).
For large $T$, the path-integral is dominated by a saddle point, leading to (2a) with
@@
(reference to $\rho_{a(b)}$ is ignored) where $(q_{\text{si}},p_{\text{si}})$ is the least-Action path for $(\rho,\widehat{\rho})$. This way, the problem reduces [31, 54] to solving the corresponding Euler-Lagrange equations
@@
where the integral in $p_{\text{si}}(x,0)$ is the contribution [31] from $\mathcal{F}$ for the initial state of the SSEP with Bernoulli measure. Additional spatial boundary conditions,
@@
are due to the “fast-coupling” [41, 58] with reservoir.
This variational problem is reminiscent of the corresponding problem on the infinite line [31] for a Bernoulli-measured initial state with an average density $\rho_{a}$ for $x<0$ and $\rho_{b}$ for $x>0$. The only differences between the two problems are the domain $x$ and the boundary conditions. The infinite line problem was formulated in [31] within MFT, which was recently solved in [39] by identifying an ingenious mapping to the classical integrable system and employing the inverse scattering method. Despite having small differences, the semi-infinite case poses a new nontrivial problem [68], which is incredibly difficult to solve.
A remarkable simplification arises for the special choice of initial density pair $(\rho_{a},\rho_{b})\equiv(1/2,0)$ for the semi-infinite line problem. For this choice of densities, there are no fluctuations in the initial state, which amounts to setting $\mathcal{F}=0$ in (11) with the condition
@@
The Euler-Lagrange equations for the corresponding variational problem (12) remains the same as in (13). The only difference comes in the initial condition in (14), which is now replaced by (16). This initial condition corresponds to the quenched averaging [31, 41].
We now show that this quenched semi-infinite line problem has a direct mapping to the quenched infinite line problem with $(\rho_{a},\rho_{b})\equiv(1,0)$ and fugacity $\tilde{\lambda}=2\lambda$. For the latter problem, the Euler-Lagrange equation is the same [31] as in (13), but now on the entire real line $x$ with the temporal boundary conditions
@@
(subscript ‘inf’ denotes infinite). It is trivial to check that the solution admits the symmetry
@@
at all times $0<t<T$. This conclusion is an essential part of our observation, as now, the solution $q_{\text{inf}}(x,t)$ on positive line $x$ satisfies the boundary conditions (15, 16) of the semi-infinite line problem with $(\rho_{a},\rho_{b})\equiv(1/2,0)$. For a similar correspondence of the response field, we define
@@
which too now replicates the boundary conditions $\tilde{p}(x,1)=\lambda$ and $\tilde{p}(0,t)=0$ of the semi-infinite line problem with $\tilde{\lambda}=2\lambda$. The fields $(q_{\text{inf}},\tilde{p})$ also satisfy the same Euler-Lagrange equations (13).
@@
for $x\geq 0$ at all times. This correspondence relates the least-Action of the two problems resulting [58] in our crucial observation:
@@
where the pre-factor $1/2$ comes from the half-domain of integration of $x$ in (11). The latter scgf is known from the seminal work [38] of Derrida and Gerschenfeld, which culminates in

### v3

# Large deviations of current for the symmetric simple exclusion process on a semi-infinite line, and on an infinite line with a slow bond

@@ Abstract
Two influential exact results in classical one-dimensional diffusive transport are about current statistics for the symmetric simple exclusion process: one in the stationary state on a finite line coupled with two unequal reservoirs at the boundaries, and the other in the non-stationary state on an infinite line. We present the corresponding result for the intermediate geometry of a semi-infinite line coupled with a single reservoir. This result is obtained using the fluctuating hydrodynamics approach of macroscopic fluctuation theory and confirmed by rare event simulations using a cloning algorithm. We apply our exact result for solving several related challenging problems, namely, the full counting statistics in presence of a defect bond, exclusion process with localized injection, survival of a tagged particle in presence of an absorbing boundary, and the stretched exponential decay in a kinetically constrained model.
large deviations, macroscopic fluctuation theory, symmetric simple exclusion process, slow bond, kinetically constrained model, stretched exponential correlation, target survival, fast injection
@@
Current fluctuations in non-equilibrium transport have long been a subject of interest in both classical and quantum contexts [2002_Prähofer_Current, 2005_Bertini_Current, *2006_Bertini_Non, 2007_Derrida_Non, 2015_Lazarescu_The, 2020_Banerjee_Current, 2023_Bello_Current, 1995_Lee_Universal, 2023_McCulloch_Full, 2024_Wienand_Emergence, turkeshi2024]. The major interest in these investigations lies in the full counting statistics in terms of large deviations of current. Apart from being a contender for extension of free energy for non-equilibrium systems, large deviation function (ldf) in general can characterize various peculiarities of non-equilibrium conditions, such as non-local response, emergent symmetries, and low-dimensional phase transitions [2005_Bodineau_Distribution, 2007_Derrida_Non, 2008_Bodineau_Long, 2010_Jona_From, 2009_Touchette_The, 2009_Bertini_Towards, 2011_Hurtado_Symmetries, 2018_Baek_Dynamical]. However, estimating ldf poses a major challenge, often requiring specialized integrability techniques [2001_Schutz_Exactly, 2015_Lazarescu_The, 2015_Mallick_The] tailored to curated models or clever numerical sampling schemes of rare events [2006_Giardinà_Direct, 2007_Lecomte_Numerical, 2011_Giardinà_Simulating, 2019_Perez_Sampling, Hartmann_computer_sim]. Understandably, exact results about ldf play an important role in the landscape of non-equilibrium physics, providing a benchmark for affirming qualitative predictions of approximate methods. Our work in this Letter presents a non-trivial addition to this list of exact results, fostering solution to further problems of practical significance [1989_Spohn_Stretched, 2012_Krapivsky_Symmetric, 2013_Franco_Phase, *2016_Franco_Phase].
@@
Among the widely studied stochastic models of classical transport, is the symmetric simple exclusion process (SSEP) [1991_Spohn_Large, 1999_Kipnis_Scaling, 1999_Liggett_Stochastic, 2007_Derrida_Non]. SSEP, along with its driven variants, has attained the status of a paradigmatic model in non-equilibrium statistical physics [2011_Chou_Non, 2015_Mallick_The]. Two celebrated exact results for SSEP concern the large deviations of current on a finite lattice coupled with two unequal reservoirs, and on an infinite lattice starting with a non-stationary state. The two geometries represent distinct non-equilibrium scenarios: a stationary state for finite systems, and a time-dependent state relaxing towards an asymptotic equilibrium state for infinite systems. They illuminate crucial differences in their fluctuations. The finite system, in the long run, holds no memory of the initial state, while the infinite system exhibits an unusual dependence on the initial state even at long times [2009_Derrida_Current2, 2012_Krapivsky_Fluctuations].
@@
The large deviations of current in these two geometries were obtained using the additivity conjecture [2004_Bodineau_Current, 2007_Bodineau_Cumulants, 2021_Derrida_Large] and integrability methods, such as the diagonalization of tilted matrix [2004_Derrida_Current] or the matrix product states [2012_Gorissen_Exact, 2015_Lazarescu_The] for the finite lattice, and the Bethe ansatz [2009_Derrida_Current] for the infinite lattice. These microscopic results were subsequently verified [2005_Bertini_Current, *2006_Bertini_Non, 2022_Mallick_Exact, *2024_Mallick_Exact, 2024_Saha_Large] using a fluctuating hydrodynamics framework [2009_Derrida_Current, 2015_Bertini_Macroscopic, 2023_Jona_Current]. This framework, now famously known as the macroscopic fluctuation theory (MFT) [2001_Bertini_Fluctuations, *2002_Bertini_Macroscopic, 2005_Bertini_Current, *2006_Bertini_Non], emerged in the early 2000s from the seminal works of Bertini, De Sole, Gabrielli, Jona-Lasinio, and Landim, which presented a general approach for characterizing non-equilibrium fluctuations of diffusive systems. MFT has successfully led to exact results of large deviations in exclusion processes [2005_Bertini_Current, *2006_Bertini_Non, 2007_Tailleur_Mapping, *2008_Tailleur_Mapping, 2015_Bertini_Macroscopic, 2022_Mallick_Exact, 2024_Saha_Large] and related transport models [2015_Bertini_Macroscopic, 2022_Bettelheim_Inverse, *2022_Bettelheim_Full, 2023_Agranov_Macroscopic, 2024_Bettelheim_Complete].
@@
In this Letter, we consider the intermediate scenario: SSEP on a semi-infinite lattice [1975_Liggett_Ergodic, 2004_Grosskinsky_Phase] coupled to a boundary reservoir with a density that is different from the initial bulk density of the system. This elucidates a non-equilibrium regime evolving into an asymptotic state in equilibrium with the reservoir [2023_Saha_Current]. In fact, the current statistics in this geometry describes the early time ($t\ll L^{2}$) statistics near the reservoir of a finite system of length $L$, extending the previous steady state ($t\gg L^{2}$) result [2004_Bodineau_Current]. The statistics have also been useful recently in the context of Mpemba effect [turkeshi2024] in quantum circuits and particle injection in a lattice [2014_Krapivsky_Lattice, 2012_Krapivsky_Symmetric]. We shall further see that the semi-infinite geometry has direct relation to transport in presence of defect and survival statistics of tracer. Despite the relevance, only limited results [2012_Williams_Combinatorics, 2013_Tracy_The, 2018_Duhart_The, 2023_Saha_Current] are known for this intermediate geometry, particularly because extending the aforementioned integrability methods for this geometry proves challenging. Even a solution via MFT has remained elusive [2023_Saha_Current].
@@
We overcame these technical challenges by uncovering a novel mapping to the infinite-line problem that leads to an exact result for the ldf of current for SSEP on the semi-infinite line, which is our main result. The mapping to the infinite-line-problem is possible due to certain symmetries of the Euler-Lagrange equations and the boundary conditions of the associated variational problem within MFT. These techniques, however, could not be obviously exploited within microscopic approaches. Our exact result is a testament to the power of MFT for non-equilibrium systems that are otherwise formidable to approach.
Our exact result for the semi-infinite SSEP enables the solution of three additional challenging problems. First, we obtain the exact ldf of current for an infinite SSEP with a localized slow bond, extending the celebrated result of Derrida and Gerschenfeld [2009_Derrida_Current]. Second, we derive the corresponding ldf for fast particle injection at a single site [2014_Krapivsky_Lattice, 2012_Krapivsky_Symmetric] in an infinite SSEP. Third, we compute the survival probability of a tracer in a fully packed SSEP with an absorbing boundary. This result extends the classical static-target [2014_Meerson_Survival] survival problem, which concerns the special case where no particle escapes the system.
@@
Model and main result: SSEP on a semi-infinite lattice (see Fig. 1), with sites indexed by $i=1,2,\cdots$, is composed of continuous-time hard-core random walkers hopping to adjacent sites with unit rate provided the target site is empty. At the boundary site $i=1$, particles are injected following exclusion with rate $\gamma\rho_{a}$ and removed at rate $\gamma(1-\rho_{a})$, which models [2007_Derrida_Non] coupling with a reservoir of density $\rho_{a}$. For each site $i$ at a given time $\tau$, we assign a binary occupation number $n_{i}(\tau)$ that takes values 0 and $1$, depending on whether the site is empty or occupied, respectively. Initially, sites are filled following Bernoulli distribution with a uniform average density $\rho_{b}$.
@@
Figure 1: SSEP on a semi-infinite lattice coupled with a reservoir of density $\rho_{a}$ with a coupling strength $\gamma>0$.
@@
Our main result concerns the time-integrated current, $Q_{T}$, which represents the total flux of particles from the reservoir into the system over a time-period $T$. In the hydrodynamic description [SM], expressed in terms of coarse-grained density $\rho(i/\sqrt{T},\tau/T)\simeq n_{i}(\tau)$, the flux
@@
$\big<\mathrm{e}^{\lambda Q_{T}}\big>\sim\mathrm{e}^{\sqrt{T}\mu_{\text{si}}(\lambda,\rho_{a},\rho_{b})}.$ (2)
The scaled cumulant generating function (scgf)
$\mu_{\text{si}}(\lambda,\rho_{a},\rho_{b})=R_{\text{si}}\big(\omega(\lambda,\rho_{a},\rho_{b})\big)$ (3a)
(subscript ‘si’ denotes semi-infinite) with
$R_{\text{si}}(\omega)=\begin{cases}R(\omega)&\text{for }\omega\geq-\frac{1}{2}\\ \displaystyle{-R(\omega)-\frac{1}{\sqrt{\pi}}\zeta\bigg(\frac{3}{2}\bigg)}&\text{for }\omega\leq-\frac{1}{2}\end{cases}$ (3b)
where $\zeta(z)$ is the Riemann zeta function and
$R\big(\omega\big)=\int_{-\infty}^{\infty}\frac{\mathrm{d}k}{2\pi}\log{\big[1+4\,\omega\,(1+\omega)\,\mathrm{e}^{-k^{2}}\big]}$ (3c)
and $\omega$ is a function of the parameters defined as [2009_Derrida_Current, 2012_Krapivsky_Fluctuations]
$\omega(\lambda,\rho_{a},\rho_{b})=\rho_{a}(1-\rho_{b})(\mathrm{e}^{\lambda}-1)+\rho_{b}(1-\rho_{a})(\mathrm{e}^{-\lambda}-1).$ (3d)
This parametric dependence of scgf on $\lambda$, $\rho_{a}$, and $\rho_{b}$ through a single function $\omega$ arises from a symmetry of the underlying dynamics of SSEP [2004_Derrida_Current, 2009_Derrida_Current, 2010_Lecomte_Current]. The piecewise function (3b) is constructed by analytic continuation of $R(\omega)$ demanding a convex scgf independently confirmed by numerical simulation (see End Matter). The expression (3) is consistent with an earlier result [2023_Saha_Current] in the low density limit.
The function $R(\omega)$ has different representations. For example, $R(\omega)=-1/(2\sqrt{\pi})\mathrm{Li}_{3/2}(-4\omega(1+\omega))$ in terms of the poly-logarithm function $\mathrm{Li}_{s}(z)$ is similar to the expression of the number of particles in Fermi-Dirac statistics. Another representation in terms of Fredholm determinant
$R(\omega)=\frac{1}{2\pi}\log{\det{[\mathbbm{1}+4\omega(1+\omega)K(x,y)]}},$ (4)
with the kernel $K(x,y)=\delta(x-y)\exp{(\partial_{x}^{2})}$, finds analogue in the current statistics for asymmetric exclusion process [2000_Johansson_Shape, 2008_Tracy_A, 2009_Tracy_Total].
It is instructive to compare (3) with the corresponding result for the infinite line [2009_Derrida_Current]. Similar to the infinite line case [2009_Derrida_Current2], the scgf (3) admits the Gallavotti-Cohen-type fluctuation symmetry [Evans1993PRL, Gallavotti1995PRL, Kurchan1998JPA, Lebowitz1999JSP]
@@
Figure 2: Scgf for semi-infinite SSEP: Solid red line represents the theoretical result of scgf (3) for $\rho_{a(b)}=0.5$, with blue dots representing the corresponding simulation result obtained by the cloning algorithm. The results closely match in the range $1\leq\lambda\leq 1$, with the deviations shown in the inset. For comparison, Gaussian approximation and the scgf for infinite line $\mu_{\text{inf}}$ [2009_Derrida_Current] are shown in dotted magenta and black dashed line, respectively.
Numerical confirmation: For confirming our result (3) based on the hydrodynamic theory for SSEP, we have independently generated the scgf using rare-event simulation based on a continuous-time cloning algorithm [SM, 2006_Giardinà_Direct, 2007_Lecomte_Numerical] with $10^{5}$ clones and measured over duration $T=250$. The data, plotted in Fig. 2, shows a good agreement with our theoretical result (3) for a reasonably large parameter range of $\lambda$. Deviations emerging at larger values of $\lambda$ are a consequence of finite-size effects [SM].
LDF: The scaling in (2) corresponds to a large deviation asymptotics of the probability
@@
where $\phi_{\text{si}}(j)$ is the ldf, related to $\mu_{\text{si}}(\lambda)$ (reference to $\rho_{a(b)}$ is ignored) by a Legendre-Fenchel transformation [2009_Touchette_The, SM] $\phi_{\text{si}}(j)=\max_{\lambda}{[j\lambda-\mu_{\text{si}}(\lambda)]}$. Using the transformation, it is immediate that (5) reflects the symmetry
@@
Similarly, the asymptotics $\mu_{\text{si}}(\lambda)\simeq[(4\sqrt{2})/(3\pi)](\log{\omega})^{3/2}$ of (3) for large positive $\lambda$ correspond [2009_Derrida_Current] to the asymptotics of the ldf
@@
Derivation: In the following, we outline our derivation of (3) within the fluctuating hydrodynamics framework of MFT. The crucial idea behind MFT is to recognize the relevant hydrodynamic modes for a coarse-grained description of the dynamics and characterize the probability of their evolution in terms of an Action, which is analogous to the Martin-Siggia-Rose-Janssen-De Dominicis (MSRJD) Action [1973_Martin_Statistical, 1976_Janssen_On, 1978_Dominicis_Dynamics, 1978_Dominicis_Field] of the associated fluctuating hydrodynamics equation. For SSEP, the relevant hydrodynamic mode is the locally conserved density $\rho(x,t)$ evolving by [2023_Saha_Current, SM]
@@
where $\sigma(\rho)=2\rho(1-\rho)$, and $\eta(x,t)$ is a delta-correlated Gaussian white noise with unit covariance. Corresponding MSRJD-Action on the semi-infinite line is [2023_Saha_Current, SM]
$S_{0}=\mathcal{F}+\int_{0}^{1}\mathrm{d}t\int_{0}^{\infty}\mathrm{d}x\bigg[\widehat{\rho}\,\partial_{t}\rho-\bigg(\frac{\sigma(\rho)}{2}\partial_{x}\widehat{\rho}+\partial_{x}\rho\bigg)\partial_{x}\widehat{\rho}\bigg$] (10)
where $\widehat{\rho}$ is the response field and $\mathcal{F}$ incorporates contributions from fluctuations in the initial state [2009_Derrida_Current2, 2023_Saha_Current]. Within this description, the generating function $\langle\mathrm{e}^{\lambda Q_{T}}\rangle=\int\mathcal{D}\left[\rho,\widehat{\rho}\right]\,\mathrm{e}^{-\sqrt{T}S_{\lambda}\left[\rho,\widehat{\rho}\right]}$ with $S_{\lambda}=S_{0}-\lambda Q_{T}/\sqrt{T}$ and $Q_{T}$ in (1).
For large $T$, the path-integral is dominated by a saddle point, leading to (2) with
@@
[…]
@@
where the integral in $p_{\text{si}}(x,0)$ is the contribution [2009_Derrida_Current2] from $\mathcal{F}$ for the initial state of SSEP with Bernoulli measure. Additional spatial boundary conditions,
@@
are due to the “fast-coupling” [2024_Saha_Large, SM] with reservoir.
This variational problem is reminiscent of the corresponding problem on the infinite line [2009_Derrida_Current2] for a Bernoulli-measured initial state with an average density $\rho_{a}$ for $x<0$ and $\rho_{b}$ for $x>0$. The only differences between the two problems are the domain $x$ and the boundary conditions. The infinite line problem was formulated in [2009_Derrida_Current2] within MFT, which was recently solved in [2022_Mallick_Exact] by identifying an ingenious mapping to the classical integrable system and employing the inverse scattering method. Despite having small differences, the semi-infinite case poses a new nontrivial problem [1975_Ablowitz_The], which is incredibly difficult to solve.
A simplification arises for the special choice of initial density pair $(\rho_{a},\rho_{b})\equiv(1/2,0)$ for the semi-infinite line problem. For this choice of densities, there are no fluctuations in the initial state, which amounts to setting $\mathcal{F}=0$ in (10) with the condition
@@
The Euler-Lagrange equations for the corresponding variational problem (11) remains the same as in (12). The only difference comes in the initial condition in (13), which is now replaced by (15). This initial condition corresponds to the quenched averaging [2009_Derrida_Current2, 2024_Saha_Large].
We now show that this quenched semi-infinite line problem has a direct mapping to the quenched infinite line problem with $(\rho_{a},\rho_{b})\equiv(1,0)$ and fugacity $\tilde{\lambda}=2\lambda$. For the latter problem, the Euler-Lagrange equation is the same [2009_Derrida_Current2] as in (12), but now on the entire real line $x$ with the temporal boundary conditions
@@
(subscript ‘inf’ denotes infinite and $\theta(x)$ is the Heaviside step function). It is simple to check that the solution admits the symmetry
@@
at all times $0<t<T$. This conclusion is an essential part of our observation, as now, the solution $q_{\text{inf}}(x,t)$ on positive line $x$ satisfies the boundary conditions (14, 15) of the semi-infinite line problem with $(\rho_{a},\rho_{b})\equiv(1/2,0)$. For a similar correspondence of the response field, we define
@@
[…]
@@
for $x\geq 0$ at all times. This correspondence relates the least-Action of the two problems resulting [SM] in our crucial observation:
@@
where the pre-factor $1/2$ comes from the half-domain of integration of $x$ in (10). The latter scgf is known from the seminal work [2009_Derrida_Current] of Derrida and Gerschenfeld, which culminates in